# Fast Estimation of the Composite Link Model for Multidimensional Grouped Counts

### v1

@@ 2.2 Use of Generalized Linear Array Models in PCLMs
Currie et al. (2006) introduced an arithmetic of arrays which enables low-storage, high-speed computations within the scoring algorithm of generalized linear models, referred to as the Generalized Linear Array Model (GLAM). In this section, we demonstrate how these principles can be effectively applied to solve (11) and estimate coefficients $\bm{\alpha}$. Each iteration of the algorithm requires only two calculations, which can be efficiently performed using the GLAM framework.
@@
$\bm{B}^{\prime}\tilde{\bm{\Gamma}}\bm{B}\equiv\rho(\mathcal{G}(\bm{B}_{2},\bm{B}_{2})^{\prime},\rho(\mathcal{G}(\bm{B}_{1},\bm{B}_{1}),\tilde{\bm{\Gamma^{*}}})).$
The symbol $\equiv$ means that both sides have the same elements but arranged in a different order, the right hand side of the equation is of size $c_{1}^{2}\times c_{2}^{2}$, therefore, it needs to be rearranged into a $c_{1}c_{2}\times c_{1}c_{2}$ matrix. Furthermore, $\mathcal{G}()$ represents the row-tensor transformation and $\rho()$ denotes the rotated $\mathcal{H}$-transform of an array, both of which were proposed by Currie et al. (2006).
@@
    $\rho(\mathcal{G}(\bm{B}_{d},\bm{B}_{d})^{\prime},\ldots,\rho(\mathcal{G}(\bm{B}_{2},\bm{B}_{2})^{\prime},\rho(\mathcal{G}(\bm{B}_{1},\bm{B}_{1})^{\prime},\tilde{\bm{\Gamma^{*}}}))\ldots).$
A detailed description of these functions, along with R code snippets, is provided in the appendix. Furthermore, a complete set of routines and fully self-reproducible programs for estimating the examples presented in this paper are accessible at osf.io/uwejt/?view_only=2ca1fdb7568342bbb9a3c51fd33c718c (temporarily anonymized to maintain confidentiality during the review process).
@@ 2.3 Uncertainty quantification
To quantify the uncertainty of the estimate $\hat{\bm{\alpha}}$, we derive its approximate distribution using a Bayesian approach proposed by Wood (2006). Under this framework:
$\bm{\alpha|y}\sim N\left(\hat{\bm{\alpha}},\bm{V}\right)\,,$ (12)
@@
$\bm{\alpha|\breve{y}}\sim N\left(\hat{\bm{\alpha}},(\bm{B}^{\prime}\bm{\Gamma}\bm{B}+\bm{P})^{-1}\right)\,.$ (13)
However, we would be underestimating the variance, since we are ignoring the uncertainty due to the redistribution of $\bm{y}$ into $\breve{\bm{y}}$. The correct expression for the information matrix can easily be obtained by calculating the derivative of (7) with respect of $\bm{\alpha}$:
@@
$\bm{\alpha|y}\sim N\left(\hat{\bm{\alpha}},(\bm{B}^{\prime}\bm{\Gamma}\bm{C}^{\prime}\bm{W}^{-1}\bm{C}\bm{\Gamma}\bm{B}^{\prime}+\bm{P})^{-1}\right).$ (15)
Furthermore, if the EM algorithm is used, the variance in (15) corresponds to the expression proposed by Lee and Pawitan (2014).
@@
    where $\bm{\Gamma}^{*}$ is a matrix of size $m_{1}\times m_{2}$ with the elements of $\hat{\bm{\gamma}}$. In this case, the matrix obtained by using GLAM methods is of size $c_{1}n_{1}\times c_{2}n_{2}$ and needs to be reshape into a matrix of size $c_{1}c_{2}\times n_{1}n_{2}$. To complete the computation of $\bm{V}$, we calculate the product $\bm{C}\bm{\Gamma}\bm{B}$, which is simply the transpose of the previously obtained $\bm{B}^{\prime}\bm{\Gamma}\bm{C}^{\prime}$. The final inner product with $\bm{W}^{-1}$ can be performed efficiently without fully constructing $\bm{W}$, as it is diagonal.
@@
    $diag"(\bm{B}\bm{V}\bm{B}^{\prime})=\verb"diag"((\bm{B}_{1}^{\prime}\otimes\bm{B}_{2}^{\prime})\bm{V}(\bm{B}_{1}\otimes\bm{B}_{2}))=\verb"vec"(\rho(\mathcal{G}(\bm{B}_{1})^{\prime},\rho(\mathcal{G}(\bm{B}_{2}),\bm{V^{*}}))),$
@@
A notable computational advantage of our approach is that the matrix $\bm{V}$ is computed only once, after the algorithm has converged. In contrast, the original estimation procedure proposed by Thompson and Baker (1981) and later adopted by Eilers (2007) requires $\bm{V}$ to be recalculated at every iteration of the algorithm, adding considerable computational overhead.
@@ 3.1 Mortality grouped by age and years
In this section, we present an illustrative example primarily aimed at demonstrating the computational efficiency of the proposed algorithm for estimating a PCLM, which has been reformulated to leverage GLAM arithmetic. Our focus here is not on evaluating the accuracy of the PCLM in recovering the latent distribution from grouped data, as this has already been demonstrated elsewhere (Rizzi et al., 2015, 2018; Lambert and Eilers, 2009; Lambert, 2011). Instead, we construct an example where the underlying distribution over single years of age and year is known. We then artificially aggregate the observed counts to simulate grouped data, allowing us to focus on computational aspects.
We use Swedish mortality data for females spanning the years 1960 to 2019, covering ages 10 to 104, i.e. $m_{1}=95$ and $m_{2}=60$. Detailed data on deaths and population exposures, provided by single year of age, is sourced from the Human Mortality Database (2024). For our analysis, we retain the original data structure for exposures while aggregating the death counts into 5-year age groups and across 5-year calendar intervals, i.e. $n_{1}=19$ and $n_{2}=12$:
@@
Figure 1 displays the actual death counts (left panel) and population exposures (right panel) as shaded contour plots across age and time, i.e. Lexis surfaces. The difference in the level of aggregation between the two datasets is clearly noticeable.
@@
Figure 2: Observed death rates by 5-year age groups and 5-year calendar intervals, along with estimated ungrouped mortality by single year of age and calendar year, for four selected age groups (15–19, 50–54, 75–79, and 95–99) over time with associated 95% confidence intervals. Data for females in Sweden, ages 10–104, from 1960 to 2019.
@@
The original procedure proposed by Thompson and Baker (1981), later adopted by Eilers (2007), and the approach proposed here for estimating a Penalized Composite Link Model both produce equivalent outcomes. However, the proposed GLAM algorithm offers a significant advantage. Its primary strength lies in its exceptional computational efficiency, making it a superior choice for large-scale applications. In the previous two-dimensional illustrative dataset, fitting the model with the proposed method and described settings takes approximately 0.75 seconds, compared to 7.7 seconds for the original algorithm (tested on a portable personal computer with an Intel i7-10610U processor, 1.8 GHz, and 16 GB of RAM). Notably, computing the variance-covariance matrices accounts for about 35% and 75% of the overall computational time for the proposed and original approach, respectively. Excluding this step, the time required for model fitting alone is reduced to about 1.6 seconds for the original method and just 0.11 seconds for the proposed approach. Furthermore, the proposed approach is highly efficient in terms of data storage. The R workspace containing all data and outcomes requires only 3.1 MB with the proposed method, whereas the conventional algorithm demands a substantially larger file size of 256 MB.

### v2

@@ 2.2 Use of Generalized Linear Array Models in PCLMs
CurrieDurbanEilersGLAM2006 introduced an arithmetic of arrays which enables low-storage, high-speed computations within the scoring algorithm of generalized linear models, referred to as the Generalized Linear Array Model (GLAM). In this section, we demonstrate how these principles can be effectively applied to solve (11) and estimate coefficients $\bm{\alpha}$. Each iteration of the algorithm requires only two calculations, which can be efficiently performed using the GLAM framework.
@@
$).\bm{B}^{\prime}\tilde{\bm{\Gamma}}\bm{B}\equiv\rho(\mathcal{G}(\bm{B}_{2},\bm{B}_{2})^{\prime},\rho(\mathcal{G}(\bm{B}_{1},\bm{B}_{1}),\tilde{\bm{\Gamma^{*}}}$)).
The symbol $\equiv$ means that both sides have the same elements but arranged in a different order, the right-hand side of the equation is of size $c_{1}^{2}\times c_{2}^{2}$, therefore, it needs to be rearranged into a $c_{1}c_{2}\times c_{1}c_{2}$ matrix. Furthermore, $\mathcal{G}()$ represents the row-tensor transformation and $\rho()$ denotes the rotated $\mathcal{H}$-transform of an array, both of which were proposed by CurrieDurbanEilersGLAM2006.
@@
[…]
A detailed description of these functions, along with R code snippets, is provided in the appendix. Furthermore, a complete set of routines and fully self-reproducible programs for estimating the examples presented in this paper are accessible at osf.io/uwejt/?view_only=2ca1fdb7568342bbb9a3c51fd33c718c.
@@ 2.3 Uncertainty quantification
To quantify the uncertainty of the estimate $\hat{\bm{\alpha}}$, we derive its approximate distribution using a Bayesian approach proposed by Wood2006. Under this framework:
$\bm{\alpha|y}\sim\mathcal{N}\left(\hat{\bm{\alpha}},\bm{V}\right)\,$, (12)
@@
$\bm{\alpha|\breve{y}}\sim\mathcal{N}\left(\hat{\bm{\alpha}},(\bm{B}^{\prime}\bm{\Gamma}\bm{B}+\bm{P})^{-1}\right)\,.$ (13)
However, we would be underestimating the variance, since we are ignoring the uncertainty due to the redistribution of $\bm{y}$ into $\breve{\bm{y}}$. The correct expression for the information matrix can easily be obtained by calculating the derivative of (7) with respect to $\bm{\alpha}$:
@@
$\bm{\alpha|y}\sim\mathcal{N}\left(\hat{\bm{\alpha}},(\bm{B}^{\prime}\bm{\Gamma}\bm{C}^{\prime}\bm{W}^{-1}\bm{C}\bm{\Gamma}\bm{B}^{\prime}+\bm{P})^{-1}\right).$ (15)
Furthermore, if the EM algorithm is used, the variance in (15) corresponds to the expression proposed by LeePawitanVarianceEM2014.
@@
    where $\bm{\Gamma}^{*}$ is a matrix of size $m_{1}\times m_{2}$ with the elements of $\hat{\bm{\gamma}}$. In this case, the matrix obtained by using GLAM methods is of size $c_{1}n_{1}\times c_{2}n_{2}$ and needs to be reshaped into a matrix of size $c_{1}c_{2}\times n_{1}n_{2}$. To complete the computation of $\bm{V}$, we calculate the product $\bm{C}\bm{\Gamma}\bm{B}$, which is simply the transpose of the previously obtained $\bm{B}^{\prime}\bm{\Gamma}\bm{C}^{\prime}$. The final inner product with $\bm{W}^{-1}$ can be performed efficiently without fully constructing $\bm{W}$, as it is diagonal.
@@
    $\verb"diag"(\bm{B}\bm{V}\bm{B}^{\prime})=\verb"diag"((\bm{B}_{1}^{\prime}\otimes\bm{B}_{2}^{\prime})\bm{V}(\bm{B}_{1}\otimes\bm{B}_{2}))=\verb"vec"(\rho(\mathcal{G}(\bm{B}_{1})^{\prime},\rho(\mathcal{G}(\bm{B}_{2}),\bm{V^{*}}))),$
@@
A notable computational advantage of our approach is that the matrix $\bm{V}$ is computed only once, after the algorithm has converged. In contrast, the original estimation procedure proposed by ThompsonBakerCLM1981 and later adopted by Eilers2007 requires $\bm{V}$ to be recalculated at every iteration of the algorithm, adding considerable computational overhead.
3 Simulation Studies
We conduct two simulation studies designed to evaluate the performance of the proposed PCLM methodology. The simulation setup is closely aligned with our motivating applications and aims to reflect realistic grouped data scenarios. Specifically, we generate high-resolution latent rates over a regular grid, apply a predefined aggregation scheme to produce grouped intensities, and simulate observed counts from a Poisson distribution. The following diagram shows the data generation process:
Design $\bm{\gamma}=\exp(\bm{\eta})$Aggregate $\bm{\mu}=\bm{C\gamma}$Simulate $\bm{y}\sim\mathcal{P}(\bm{\mu})$
In the first subsection, we replicate this process multiple times within a two-dimensional setting under various underlying scenarios to evaluate the performance of the PCLM. In Section 3.2, we extend the analysis to three- and four-dimensional settings to demonstrate the computational advantages of the proposed approach.
3.1 Assessing the performance of the Penalized Composite Link Model
In this section, we evaluate the performance of the PCLM in ungrouping Poisson-distributed data within a two-dimensional framework, where observations are aggregated along both dimensions at varying levels of grouping that mimic real-world data applications. Similar evaluations have been conducted in previous studies for one-dimensional (RizziEtAlPCLM1D2015; LambertEilersBayesGrouped2009) and two-dimensional (LambertBayesBivariateCSDA2011) settings.
To conduct a comprehensive assessment under controlled yet realistic conditions, we devised two simulation scenarios, labeled A and B, each defined by a distinct true underlying linear predictor $\bm{\eta}$. In both scenarios, data are simulated over a two-dimensional covariate grid defined by $\bm{x}_{1}=1,\dots,m_{1}=80$ and $\bm{x}_{2}=1,\dots,m_{2}=60$, resulting in a total of $m_{1}\times m_{2}=4{,}800$ latent observations.
Scenario A uses a synthetic specification of $\bm{\eta}$ designed to reflect plausible nonlinear and interaction structures. Specifically, we define two coefficient functions over the temporal dimension $\bm{x}_{2}$: $\bm{\beta}_{1}=-10+0.5\cos(\bm{x}_{2}/40),\quad\bm{\beta}_{2}=0.1+0.025\cos(\bm{x}_{2}/40)$. And we set a fixed vector $\bm{\beta}_{3}=\bm{1}_{n_{2}}$. The covariate information is represented by the design matrix $\bm{X}=\left[\bm{1}_{m_{1}}:\bm{x}_{1}:-\sin\left(\frac{\pi\bm{x}_{1}}{50}\right)\right]$. This setting corresponds to a varying-coefficient model in which both the intercept and slope vary smoothly with $\bm{x}_{2}$. It is designed to emulate real-world mortality patterns, where rates change gradually and non-linearly across age and time. The structure allows for a controlled and interpretable representation, with linear predictor values ranging approximately from 10-10 to 1.5.
Scenario B, in contrast, relies on empirically observed demographic data: we use age-specific mortality rates for Swedish females from age 20 to 99 ($\bm{x}_{1}$) and years 1960–2019 ($\bm{x}_{2}$), applying light $P$-spline smoothing to obtain a realistic underlying linear predictor. The dimensional structure in this scenario matches that of Scenario A, with $\bm{x}_{1}$ and $\bm{x}_{2}$ forming an $80\times 60$ grid. In this case, the smoothed linear predictor exhibits values ranging from 8.4-8.4 to 0.7-0.7, reflecting typical mortality patterns on the log scale.
In both cases, the predictor $\bm{\eta}$ is used to compute a vector of latent distribution $\bm{\gamma}$, and expected Poisson counts are obtained by incorporating an exposure matrix $\bm{e}$, which varies across both covariate dimensions. Although exposures simply act as an offset in the model, we design $\bm{e}$ to emulate real populations, with decreasing exposure by $\bm{x}_{1}$ and smoothly varying patterns over $\bm{x}_{2}$. Two exposure settings (small and large) are considered, differing by a factor of 20 in magnitude. To contextualize the large exposure setting, it corresponds to population sizes ranging approximately from 15 to 30 million. This setup yields four distinct configurations, defined by the combination of two linear predictor scenarios (A and B) with the two exposure levels (small and large).
To assess the robustness of the PCLM under data aggregation, we simulate counts from the latent Poisson distribution and apply grouping along both dimensions using interval widths of 1, 2, 5, and 10, resulting in 16 distinct grouping schemes, including the fully ungrouped case. For illustration, the coarsest grouping—using interval widths of 10 along both dimensions—produces only $n_{1}=8$ and $n_{2}=6$ observed cells, totaling 48 aggregated data points. In contrast, grouping with interval widths of 5 yields 192 observed values ($n_{1}=16$, $n_{2}=12$). Across all configurations, the goal remains the same: to recover the original latent distribution defined over the full grid of size $m_{1}\times m_{2}=4{,}800$.
For each of the 64 (4 $\times$ 16) setting combinations, we apply the PCLM, use rich $B$-spline bases over the two dimensions ($c_{1}=16$ and $c_{2}=12$) and optimize the smoothing parameters via minimization of the Bayesian Information Criterion (Schwarz1978). This enables estimation of the underlying $\bm{\eta}$ across various grouping structures. Each configuration is replicated 100 times, and we compute the Root Mean Square Error (RMSE) between the estimated and true $\bm{\eta}$ to assess reconstruction accuracy.
Figure 1: Root Mean Square Error (RMSE) of the estimated linear predictor relative to the true underlying values, based on 100 simulation replicates for each setting in a two-dimensional framework. Boxplots show the distribution of RMSE across different grouping schemes, defined by combinations of interval widths along the two dimensions, which correspond to different numbers of observed data points. Each panel represents one of the four combinations of linear predictor scenarios (A and B) and exposure levels (Small and Large).
Outcomes are shown in Figure 1 across the 16 different grouping schemes and four simulation settings. Each boxplot summarizes the distribution of RMSE over 100 replicates for a given grouping configuration, denoted by the interval widths over the two dimensions $\bm{x}_{1}$ and $\bm{x}_{2}$. Overall, across all simulation settings, RMSE values remain small, even under substantial data aggregation, indicating that the PCLM performs effectively in ungrouping and smoothing Poisson-distributed data observed in grouped form. Numerically, RMSE values range from approximately 0.01 to 0.05. Given that the true linear predictor in both scenarios spans a range from 10-10 to 1.5, these errors correspond to less than 0.5% of the total scale.
However, a clear, albeit more modest than anticipated, trend emerges: as the level of aggregation increases (e.g., interval widths of 5 or 10), estimation error tends to rise. However, even with coarser groupings, the resulting RMSE values remain moderate, indicating that the PCLM retains robustness under substantial aggregation. In contrast, finer groupings (e.g., 1–2, 2–1) consistently yield lower RMSE values, particularly under the large exposure setting, approaching the accuracy observed with fully disaggregated data (1–1). The influence of exposure level is more pronounced: for both scenarios, RMSE values are systematically lower and exhibit reduced variability when exposure is large, reflecting the stabilizing effect of higher expected counts on estimation accuracy.
Scenario A, which is based on analytically constructed data, generally yields slightly lower RMSE values compared to Scenario B. This likely reflects the more regular and controlled structure of the underlying linear predictor in Scenario A. Overall, these results confirm that the PCLM is robust in recovering the true linear predictor, even under substantial data aggregation, particularly when exposure levels are high and grouping intervals are not excessively coarse.
3.2 Comparing computational performances
To assess the computational efficiency of our proposed GLAM-based implementation of the PCLM, we conducted a dedicated simulation study focusing on differences in storage and runtime performance between the original algorithm and the GLAM-based alternative. In contrast to previous simulations, where the emphasis was on estimation accuracy, the goal here is to evaluate how each implementation scales computationally when applied to higher-dimensional data structures.
We considered data structures of increasing dimensionality, ranging from two to four dimensions, each defined over regularly spaced covariate grids. To reflect more realistic scenarios, we examined a relatively large data configuration where $\bm{x}_{1}=\bm{x}_{2}=1,\dots,50$ and $\bm{x}_{3}=\bm{x}_{4}=1,\dots,20$, i.e., $m_{1}=m_{2}=50$ and $m_{3}=m_{4}=20$, resulting in a latent vector $\bm{\gamma}$ of length 2,500, 50,000, and 1,000,000 in the two-, three-, and four-dimensional cases, respectively. For comparison purposes, we also considered a smaller, simplified example with $m_{1}=m_{2}=40$ and $m_{3}=m_{4}=8$, for which the corresponding lengths of $\bm{\gamma}$ are 1,600, 12,800, and 102,400.
Over each of these domains and both data structures, we constructed smooth univariate functions as follows: $\bm{f}_{1}=(\sin(\bm{x}_{1}/20)+1)/2$, $\bm{f}_{2}=-4(\cos(\bm{x}_{2}/20)+1)$, $\bm{f}_{3}=\sin(\bm{x}_{3}/30)$, and $\bm{f}_{4}=\cos(\bm{x}_{4}/40)$. The true underlying linear predictor was constructed in each setting by applying successive outer products across dimensions, starting from the two-dimensional case (using $\bm{f}_{1}$ and $\bm{f}_{2}$) and progressively incorporating $\bm{f}_{3}$, and $\bm{f}_{4}$ for the three-, and four-dimensional settings, respectively.
From the resulting linear predictor arrays, Poisson-distributed count data were simulated. To maintain comparability with earlier simulations and future applications, expected values for the Poisson distribution were defined using both the exponential of the linear predictor and an exposure term. In this setting, exposures were kept constant at 10,00010{,}000 across the entire covariate grid as a simplifying assumption. For all configurations, we applied grouping only along the first two dimensions ($\bm{x}_{1}$ and $\bm{x}_{2}$) by aggregating every five consecutive units, thereby introducing coarsening that mirrors real-world data aggregation practices. Specifically, the resulting dimensions are $n_{1}=n_{2}=12$ in the realistic scenario, and $n_{1}=n_{2}=8$ in the smaller scenario.
To isolate differences in computational performance, we fixed the model complexity across all dimensions by selecting a relatively rich tensor-product $B$-spline basis with internal knots every 5 available data-points along each corresponding dimension. For simplicity and consistency, smoothing parameters were fixed at a subjectively chosen value of 100 across all dimensions. This setting allows a fair and focused comparison of memory usage and computation time between the standard and GLAM-based implementations of the PCLM across varying dimensional complexities. While computer performance was not explicitly tested in previous simulation studies, all models for this study, as well as for the applications presented in Section 4, were run on a portable personal computer equipped with an Intel i7-10610U processor (1.8 GHz) and 16 GB of RAM. This detail is important for interpreting the observed memory demands and computational times.
Memory usage (MB) Computation time (s)
Data
Dimension
(length of $\bm{\gamma}$)
Algorithm:
All
objects
w/o variance
estimation
Complete
estimation
w/o variance
estimation
[…]
Original 22.47 1.95 0.26 0.05
[…]
Original 1394.76 82.89 43.77 7.16
[…]
Original OOM 4290.90 1894.46
[…]
Original 113.39 9.52 1.55 0.35
[…]
Original OOM 2022.93 483.19
[…]
Original OOM OOM
Table 1: Memory usage (MB) and computation time (s) for estimating a Penalized Composite Link Model across two-, three-, and four-dimensional data structures. Results compare the proposed GLAM-based algorithm with the original estimation procedure. The label OOM indicates an Out Of Memory error, meaning the procedure failed due to insufficient memory on the test machine. The symbol ✗ denotes that the corresponding value could not be computed for this reason.
The results in Table 1 clearly highlight the computational advantages of the proposed GLAM-based algorithm over the original estimation procedure for Penalized Composite Link Models (PCLM) across varying data dimensions and sizes. In all scenarios considered, the GLAM-based approach consistently demonstrates substantially lower memory usage and faster computation times. These improvements become more pronounced as the dimensionality and data size increase. For instance, in the small 3D case ($|\bm{\gamma}|=12{,}800$), the original algorithm requires approximately 1,395 MB of memory for full estimation, whereas the GLAM-based algorithm completes the same task using just 7.3 MB, yielding a nearly 190-fold reduction. Likewise, computation time drops from 43.77 seconds to 0.36 seconds.
For relatively large settings, while the original algorithm is sometimes able to estimate the parameter vector $\bm{\gamma}$, it often fails to compute the associated variance-covariance matrix due to excessive memory requirements. This limitation prevents quantification of the uncertainty in the estimates, which is critical for inference and practical applications. For example, in the small 4D case ($|\bm{\gamma}|=102{,}400$), the GLAM-based method completes full estimation—including variance estimation—in approximately 13.81 seconds, whereas the original approach requires over 31 minutes (1894.46 seconds) just to estimate the point estimates without successfully computing the variance-covariance matrix.
The benefits of the GLAM-based approach are even more apparent in higher-dimensional scenarios. In the large 4D setting ($|\bm{\gamma}|=1{,}000{,}000$), the original algorithm is unable to proceed at all, whereas the proposed method, while demanding in absolute terms (around 2.6 GB and 525 seconds), remains fully operational and scalable.
Overall, the GLAM-based algorithm markedly improves the feasibility of fitting PCLMs in high-dimensional settings. It allows users to perform comprehensive estimation tasks on large and complex datasets that would otherwise exceed memory limitations with the original method.
In terms of memory usage, the primary bottleneck of the original algorithm lies in the construction of the full model and composite matrices. For example, in a relatively small four-dimensional scenario, the model matrix $\bm{B}=\bm{B}_{4}\otimes\bm{B}_{3}\otimes\bm{B}_{2}\otimes\bm{B}_{1}$ and the composite matrix $\bm{C}=\mathrm{diag}(m_{4})\otimes\mathrm{diag}(m_{3})\otimes\bm{C}_{2}\otimes\bm{C}_{1}$ require approximately 800 MB and 3,200 MB of memory, respectively. Consequently, even before any computation, the storage of $\bm{B}$ and $\bm{C}$ represents a significant constraint, especially in resource-limited environments such as personal computers.
These results highlight not only the efficiency but also the robustness of the proposed method for practical applications involving multi-dimensional smoothing and large-scale data where count data are grouped, and the objective is to estimate latent underlying distributions.
@@ 4.1 Mortality grouped by age and years
In this section, we present an illustrative example primarily aimed at demonstrating the computational efficiency of the proposed algorithm for estimating a PCLM, which has been reformulated to leverage GLAM arithmetic. We construct an example where the underlying distribution over single years of age and year is known. We then artificially aggregate the observed counts to simulate grouped data, allowing us to focus on computational aspects.
We use ungrouped mortality data for Swedish females obtained from the HMD, spanning the years 1960 to 2019 and ages 10 to 104. The data are originally reported at a fine resolution, with death counts and population exposures recorded by single year of age and single calendar year. This structure defines a high-resolution grid of size $m_{1}=95$ (ages) by $m_{2}=60$ (years), yielding a total of 5,700 latent data points, which are later smoothed using a standard two-dimensional $P$-spline approach for illustrative purposes.
For our analysis, we aggregate the original death counts into 5-year age groups and 5-year calendar intervals. This results in an observed dataset with $n_{1}=19$ age groups and $n_{2}=12$ time intervals, for a total of $n_{1}\times n_{2}=228$ observed cells:
@@
Note that the population exposures were retained at their original fine scale and used as offsets in the model, consistent with the common availability of detailed population denominators even when mortality data are aggregated. Figure 2 displays the actual death counts (left panel) and population exposures (right panel) as shaded contour plots across age and time, i.e. Lexis surfaces. The difference in the level of aggregation between the two datasets is clearly noticeable.
@@
Figure 3: Observed death rates by 5-year age groups and 5-year calendar intervals, along with estimated ungrouped mortality by single year of age and calendar year, for four selected age groups (15–19, 50–54, 75–79, and 95–99) over time with associated 95% confidence intervals. Data for females in Sweden, ages 10–104, from 1960 to 2019. For illustrative purposes, we also include (in dashed lines) the estimates that would have been obtained using a standard two-dimensional $P$-splines approach applied directly to the original unaggregated data.
@@
For illustrative purposes, we additionally apply a standard two-dimensional $P$-spline smoothing technique directly to the original ungrouped mortality data. The resulting estimates are shown as dashed lines in Figure 3. This smoothed surface, obtained from the full-resolution data (5,700 data points), serves as a reference benchmark. It enables a visual and quantitative comparison to the estimates produced by the proposed PCLM when fitted to substantially more coarsely grouped data (228 aggregated observations). This comparison helps assess the model’s ability to recover the essential structure and temporal evolution of the underlying mortality surface, even in the presence of considerable information loss due to aggregation.
To estimate the PCLM on the coarsely grouped data, we implement both the original iterative procedure introduced by ThompsonBakerCLM1981 and later extended by Eilers2007, as well as the computationally efficient approach proposed in this paper. While both methods produce equivalent results, thereby confirming the correctness of our implementation, the proposed GLAM algorithm offers a significant advantage. Its primary strength lies in its exceptional computational efficiency, making it a superior choice for large-scale applications. In the previous two-dimensional illustrative dataset, fitting the model with the proposed method and described settings takes approximately 0.14 seconds, compared to 7.9 seconds for the original algorithm. Notably, computing the variance-covariance matrices accounts for about 24% and 79% of the overall computational time for the proposed and original approach, respectively. Excluding this step, the time required for model fitting alone is reduced to about 1.7 seconds for the original method and just 0.11 seconds for the proposed approach. Furthermore, the proposed approach is highly efficient in terms of data storage. With the proposed method, the combined size of all relevant R objects used for data and estimation is only 4.4 MB, compared to 284 MB required by the conventional algorithm, a substantial reduction in memory usage.